# Learning Representations for Independence Testing

### v2

@@ 1 Introduction
In both cases, the overall approach builds on prior work in kernel two-sample testing (Gretton et al., 2012b; Jitkrittum et al., 2016; Sutherland et al., 2017; Liu et al., 2020). We show, however, that performing the direct reduction to two-sample testing and applying these techniques, in addition to breaking the assumptions of their theoretical results, yields a notably less powerful test than our direct HSIC power optimization, due to statistical properties of the estimators. We also note that Ren et al. (2024) recently proposed a closely related scheme for the HSIC power optimization setting; unfortunately, however, we show that several of their main algorithmic suggestions are misguided and the proof of their main theorem is incorrect.
@@
These are baselines considered in this paper. See more details in Section 6.
@@ 2.1 Asymptotic test power
Taking this into account, we examine the behavior of a permutation test based on (1) or the many other mutual information bounds based only on the NDS (2). Each bound will choose a different $f$ during training, but at test time only the NDS matters.
@@
Let $T_{\mathfrak{H}_{1}}=\operatorname*{\mathbb{E}}_{\mathbb{P}_{xy}}f(x,y)$ and $T_{\mathfrak{H}_{0}}=\operatorname*{\mathbb{E}}_{\mathbb{P}_{x}\times\mathbb{P}_{y}}f(x,y)$. Assuming $0<\tau_{\mathfrak{H}_{1}}^{2}=\operatorname*{Var}_{\mathbb{P}_{xy}}f(x,y)<\infty$, the central limit theorem implies that $\frac{1}{\tau_{\mathfrak{H}_{1}}}\sqrt{m}(\hat{T}-T_{\mathfrak{H}_{1}})\stackrel{{\scriptstyle\mathrm{d}}}{{\to}}\mathcal{N}(0,1)$. Then the rejection threshold $r_{m}$ should be such that, denoting $a\sim b$ to mean $\lim_{m\to\infty}a/b=1$ and using $\Phi$ for the standard normal cdf,
$\alpha=\Pr_{\mathfrak{H}_{0}}(\hat{T}>r_{m})=\Pr_{\mathfrak{H}_{0}}\biggl{(}\frac{\hat{T}-T_{\mathfrak{H}_{0}}}{\tau_{\mathfrak{H}_{0}}/\sqrt{m}}>\frac{r_{m}-T_{\mathfrak{H}_{0}}}{\tau_{\mathfrak{H}_{0}}/\sqrt{m}}\biggr{)}\sim\Phi\left(\frac{T_{\mathfrak{H}_{0}}-r_{m}}{\tau_{\mathfrak{H}_{0}}/\sqrt{m}}\right),$
and hence $r_{m}\sim T_{\mathfrak{H}_{0}}-\Phi^{-1}(\alpha){\tau_{\mathfrak{H}_{0}}}/{\sqrt{m}}$. This asymptotic threshold, combined with the asymptotic normality of $\hat{T}$ under $\mathfrak{H}_{1}$, tells us the power satisfies (recalling $T_{\mathfrak{H}_{1}}$, $\tau_{\mathfrak{H}_{1}}$, $T_{\mathfrak{H}_{0}}$, $\tau_{\mathfrak{H}_{0}}$ do not depend on $m$)
@@
$\Pr_{\mathfrak{H}_{1}}\left(\hat{T}>r_{m}\right)=\Pr_{\mathfrak{H}_{1}}\left(\frac{\hat{T}-T_{\mathfrak{H}_{1}}}{\tau_{\mathfrak{H}_{1}}/\sqrt{m}}>\frac{r_{m}-T_{\mathfrak{H}_{1}}}{\tau_{\mathfrak{H}_{1}}/\sqrt{m}}\right)\sim\Phi\left(\frac{T_{\mathfrak{H}_{1}}-r_{m}}{\tau_{\mathfrak{H}_{1}}/\sqrt{m}}\right)\sim\Phi\left(\sqrt{m}\,\frac{T_{\mathfrak{H}_{1}}-T_{\mathfrak{H}_{0}}}{\tau_{\mathfrak{H}_{1}}}+\frac{\tau_{\mathfrak{H}_{0}}}{\tau_{\mathfrak{H}_{1}}}\Phi^{-1}(\alpha)\right).$ (3)
@@
To confirm that the asymptotics here give a reasonable approximation for finite sample sizes, we can check that the asymptotic test power (3) for a given critic function $f$ roughly agrees with the empirical test power using a non-asymptotic rejection threshold estimated by obtaining fresh samples from the null distribution, i.e. the empirical quantile of $\hat{T}$ under the null. (This would not be available in practical situations, but is in synthetic problems where we can obtain arbitrary numbers of fresh samples.) Figure 2 indeed confirms that (3) gives a good estimate of the empirical power even at moderate sample sizes. (Intriguingly, the power of a permutation test is consistently slightly higher than the test based on the correct non-asymptotic threshold; we will discuss this issue in Section 7.)
@@
This argument shows that, as long as $T_{\mathfrak{H}_{1}}>T_{\mathfrak{H}_{0}}$ and $\hat{T}$ has finite positive variance under both $\mathbb{P}_{xy}$ and $\mathbb{P}_{x}\times\mathbb{P}_{y}$, a test based on $\hat{T}$ will almost surely eventually reject any fixed alternative: the test is consistent. How quickly in $m$ it reaches high power, however, depends on the expression (3), which in particular for large $m$ is dominated by the signal-to-noise ratio (SNR) given by $(T_{\mathfrak{H}_{1}}-T_{\mathfrak{H}_{0}})/\tau_{\mathfrak{H}_{1}}$. When we choose a test based on maximizing the value of a mutual information lower bound, we maximize a criterion, e.g., (1), which does not directly correspond to the test power. We will discuss instead maximizing the asymptotic power in Section 5.
@@
[…]
@@ 3 Tests with the Hilbert-Schmidt Independence Criterion (HSIC)
Suppose we have a kernel $k$ on $\mathcal{X}$ with RKHS $\mathcal{F}$ and feature map $\phi$, as well as another kernel $l$ on $\mathcal{Y}$ with RKHS $\mathcal{G}$ and feature map $\psi$. Let $\otimes$ denote the outer product. [^1] The Euclidean outer product $ab^{\mathsf{T}}$ is a matrix with $[ab^{\mathsf{T}}]b^{\prime}=a[b^{\mathsf{T}}b^{\prime}]$; in Hilbert spaces, $f\otimes g:\mathcal{G}\to\mathcal{F}$ has $[f\otimes g]g^{\prime}=f\langle g,g^{\prime}\rangle_{\mathcal{G}}$. The cross-covariance operator is
@@ 3.1 Asymptotic test power
The unbiased HSIC estimator is asymptotically normal with $\sqrt{m}(\widehat{\operatorname{HSIC}}_{\rm u}-\operatorname{HSIC})\overset{\mathrm{d}}{\to}\mathcal{N}(0,\sigma^{2}_{\mathfrak{H}_{1}})$ (Proposition A.1, in Appendix A). It is possible, however, for $\sigma^{2}_{\mathfrak{H}_{1}}=0$, which makes that result less useful; this occurs in particular whenever $\operatorname{HSIC}=0$, and hence is always true under $\mathfrak{H}_{0}$. In that case, it is more informative to say that instead $m\,\widehat{\operatorname{HSIC}}_{\rm u}$ converges in distribution (which is not the case when $\sigma^{2}_{\mathfrak{H}_{1}}=0$). The distribution to which it converges is a mixture of shifted chi-squareds with complex dependence on $\mathbb{P}_{x}$, $\mathbb{P}_{y}$, $k$, and $l$ (Proposition A.2). Thus, if we consider a test statistic $m\,\widehat{\operatorname{HSIC}}_{\rm u}$, when $\operatorname{HSIC}=0$ this statistic has mean zero and standard deviation $\Theta(1)$. When $\operatorname{HSIC}>0$, though, the statistic has mean and standard deviation $\Theta(\sqrt{m})\to\infty$. Thus, as $m\to\infty$, eventually the test will reject if $\operatorname{HSIC}>0$ (Rindt et al., 2021).
@@
Recall that the optimal test is one with highest test power among those with appropriate level control. Like for NDS-based tests, we can describe this power asymptotically. Let $\Phi$ denote the standard normal cdf, and $\Psi$ be the cdf of the null distribution of $m\,\widehat{\operatorname{HSIC}}_{\rm u}$ (which depends on kernels $k$, $l$, and distributions $\mathbb{P}_{x}$ and $\mathbb{P}_{y}$). It follows that the rejection threshold $r$ should satisfy
@@
$\alpha=\Pr_{\mathfrak{H}_{0}}(\widehat{\operatorname{HSIC}}_{\rm u}>r_{m})=\Pr_{\mathfrak{H}_{0}}(m\,\widehat{\operatorname{HSIC}}_{\rm u}>m\,r_{m})\sim 1-\Psi(m\,r_{m}),$
@@
implying that $r_{m}\sim\frac{1}{m}\Psi^{-1}(1-\alpha)$. This threshold, combined with Proposition A.1, tells us that when $\sigma_{\mathfrak{H}_{1}}>0$ the asymptotic test power satisfies
@@
$\Pr_{\mathfrak{H}_{1}}(\widehat{\operatorname{HSIC}}_{\rm u}>r_{m})=\Pr_{\mathfrak{H}_{1}}\left(\frac{\widehat{\operatorname{HSIC}}_{\rm u}-\operatorname{HSIC}}{\sigma_{\mathfrak{H}_{1}}/\sqrt{m}}>\frac{r_{m}-\operatorname{HSIC}}{\sigma_{\mathfrak{H}_{1}}/\sqrt{m}}\right)\sim\Phi\left(\frac{\sqrt{m}\operatorname{HSIC}}{\sigma_{\mathfrak{H}_{1}}}-\frac{\Psi^{-1}(1-\alpha)}{\sqrt{m}\sigma_{\mathfrak{H}_{1}}}\right).$ (6)
@@
Figure 3 verifies that this expression lines up well with the empirical test power even at moderate test sizes. Additionally, (6) also tells us that, as long as $\operatorname{HSIC}>0$ and $\widehat{\operatorname{HSIC}}_{\rm u}$ has finite positive variance, an HSIC-based test will eventually reject with probability of one. The rate at which it increases in power is described by the gap between the signal-to-noise ratio $\operatorname{HSIC}/\sigma_{\mathfrak{H}_{1}}$ and a threshold-dependent term $\Psi^{-1}(1-\alpha)/\sigma_{\mathfrak{H}_{1}}$, with the former dominating the latter. This gap increases at a faster rate than the one described for NDS-based tests (3); the ratio of the two terms is $\Theta(m)$ for HSIC, while it is only $\Theta(\sqrt{m})$ for NDS.
@@
[…]
@@ 3.2 MMD-based independence test
If we only have access to paired samples $\mathbb{Z}=\bigl{(}(x_{1},y_{1}),\dots,(x_{m},y_{m})\bigr{)}$ drawn iid from the joint distribution, we can simulate the null distribution $\mathbb{P}_{x}\times\mathbb{P}_{y}$ by shuffling the samples $\mathbb{Y}=(y_{1},...,y_{m})$ by some permutation $\sigma:[m]\to[m]$. We define the action of permuting the samples $\mathbb{Z}$ with $\sigma$ as $\sigma\mathbb{Z}=\bigl{(}(x_{1},y_{\sigma(1)}),\dots,(x_{m},y_{\sigma(m)})\bigr{)}$. One simple way to estimate $\operatorname{MMD}^{2}(\mathbb{P}_{xy},\mathbb{P}_{x}\times\mathbb{P}_{y})$ is then simply $\widehat{\operatorname{MMD}}_{b}(\mathbb{Z},\sigma\mathbb{Z})$.
@@
This result is proved in Section E.4. Since under $\mathfrak{H}_{0}$ the distribution of $\mathbb{Z}$ agrees with that of $\sigma\mathbb{Z}$ for any choice of $\sigma$, Theorem 2 of Hemerik & Goeman (2018) guarantees that any permutation test on samples $\mathbb{Z}$ against $\sigma\mathbb{Z}$ will have exactly valid level, provided the original sample order is included in the permutations. With a sufficient kernel, this estimator thus yields an exactly valid, consistent test.
One might be inclined to split the data, using $\widehat{\operatorname{MMD}}_{b}(\mathbb{Z}_{1},\sigma\mathbb{Z}_{2})$; our analysis shows this is not needed either for level control or for consistency, and indeed we find it performs worse than not splitting in practice.
We can also consider using multiple permutations, rather than just one: $\boldsymbol{\sigma}\mathbb{Z}=\bigcup\sigma_{i}\mathbb{Z}$ for permutations $\sigma_{1},\sigma_{2},\dots$. In fact, $\widehat{\operatorname{HSIC}}$ is exactly this estimator when considering all possible permutations; we show in Proposition E.2 that doing so improves the estimation quality.
So for any given dataset, the single sample of random $\sigma$ is closer and closer to its mean which is closer and closer to the true expectation, so consistent.

### v3

@@ 1 Introduction
In both cases, the overall approach builds on prior work in kernel two-sample testing (Gretton et al., 2012b; Jitkrittum et al., 2016; Sutherland et al., 2017; Liu et al., 2020). We show, however, that performing the direct reduction to two-sample testing and applying these techniques, in addition to breaking the assumptions of their theoretical results, yields a notably less powerful test than our direct HSIC power optimization, due to statistical properties of the estimators. We also note that Ren et al. (2024) recently proposed a closely related scheme for the HSIC power optimization setting; we argue, however, that several of their main algorithmic suggestions are not well-supported (particularly in Section˜F.5) and identify an apparent gap in the proof of their main theorem (see Footnote˜10).
@@
[…]
@@ 2.1 Asymptotic test power
Let $T_{\mathfrak{H}_{0}}=\operatorname*{\mathbb{E}}_{\mathbb{P}_{x}\times\mathbb{P}_{y}}f(x,y)$ and $T_{\mathfrak{H}_{1}}=\operatorname*{\mathbb{E}}_{\mathbb{P}_{xy}}f(x,y)$ be the population level statistics under the null and alternative distributions. Assuming a fixed critic $f$ is chosen independently from test samples $\mathbb{X}$ and $\mathbb{Y}$ (e.g. because of sample splitting), and that $0<\tau_{\mathfrak{H}_{1}}^{2}=\operatorname*{Var}_{\mathbb{P}_{xy}}f(x,y)<\infty$, the central limit theorem implies that $\frac{1}{\tau_{\mathfrak{H}_{1}}}\sqrt{m}(\hat{T}-T_{\mathfrak{H}_{1}})\stackrel{{\scriptstyle\mathrm{d}}}{{\to}}\mathcal{N}(0,1)$. Then, using $\Phi$ for the standard normal cdf, the rejection threshold $r_{m}$ satisfies
$\alpha=\Pr_{\mathfrak{H}_{0}}\left(\sqrt{m}\hat{T}>r_{m}\right)=\Pr_{\mathfrak{H}_{0}}\left(\sqrt{m}\frac{\hat{T}-T_{\mathfrak{H}_{0}}}{\tau_{\mathfrak{H}_{0}}}>\frac{r_{m}-\sqrt{m}T_{\mathfrak{H}_{0}}}{\tau_{\mathfrak{H}_{0}}}\right)=1-\Phi\left(\frac{r_{m}-\sqrt{m}T_{\mathfrak{H}_{0}}}{\tau_{\mathfrak{H}_{0}}}\right)+o(1).$
As $\Phi$ is Lipschitz, this implies $r_{m}=\sqrt{m}T_{\mathfrak{H}_{0}}+\tau_{\mathfrak{H}_{0}}\bigl(\Phi^{-1}(1-\alpha)+o(1)\bigr)$. Using the asymptotic normality of $\hat{T}$ under $\mathfrak{H}_{1}$, we then have that
$\Pr_{\mathfrak{H}_{1}}\left(\sqrt{m}\hat{T}>r_{m}\right)$ $=\Pr_{\mathfrak{H}_{1}}\left(\sqrt{m}\frac{\hat{T}-T_{\mathfrak{H}_{1}}}{\tau_{\mathfrak{H}_{1}}}>\frac{r_{m}-\sqrt{m}T_{\mathfrak{H}_{1}}}{\tau_{\mathfrak{H}_{1}}}\right)$
@@
$=1-\Phi\left(\frac{\sqrt{m}T_{\mathfrak{H}_{0}}+\tau_{\mathfrak{H}_{0}}\bigl(\Phi^{-1}(1-\alpha)+o(1)\bigr)-\sqrt{m}T_{\mathfrak{H}_{1}}}{\tau_{\mathfrak{H}_{1}}}\right)+o(1)$
$=\Phi\left(\sqrt{m}\frac{T_{\mathfrak{H}_{1}}-T_{\mathfrak{H}_{0}}}{\tau_{\mathfrak{H}_{1}}}-\frac{\tau_{\mathfrak{H}_{1}}}{\tau_{\mathfrak{H}_{0}}}\bigl(\Phi^{-1}(1-\alpha)+o(1)\bigr)\right)+o(1).$ (3)
@@
To confirm that these asymptotics give a reasonable approximation in finite sample sizes, we can check that the asymptotic test power (3) for a given critic function $f$ roughly agrees with the empirical test power using a non-asymptotic rejection threshold estimated by obtaining fresh samples from the null distribution, i.e. the empirical quantile of $\hat{T}$ under the null. (This would not be available in practical situations, but is in synthetic problems where we can obtain arbitrary numbers of fresh samples.) Figure˜2 indeed confirms that (3) gives a good estimate of the empirical power even at moderate sample sizes. (Intriguingly, the power of a permutation test is consistently slightly higher than the test based on the correct non-asymptotic threshold; we will discuss this issue in Section˜7.)
@@
This argument shows that, as long as $T_{\mathfrak{H}_{1}}>T_{\mathfrak{H}_{0}}$ and $\hat{T}$ has finite positive variance under both $\mathbb{P}_{xy}$ and $\mathbb{P}_{x}\times\mathbb{P}_{y}$, a test based on $\hat{T}$ will almost surely eventually reject any fixed alternative: the test is consistent. How quickly in $m$ it reaches high power, however, depends on the expression (3), which in particular for large $m$ is dominated by the signal-to-noise ratio (SNR) given by $(T_{\mathfrak{H}_{1}}-T_{\mathfrak{H}_{0}})/\tau_{\mathfrak{H}_{1}}$. When we choose a test based on maximizing the value of a mutual information lower bound, we maximize a criterion, e.g., (1), which does not directly correspond to the test power. We will discuss instead maximizing the asymptotic power in Section˜5.
@@
Figure 2: Power of NDS tests with a particular critic $f$ (learned by the approach of Section˜5) on three datasets described in Section˜6. For each problem, the blue line (asymp) is power described by the asymptotic formula (3) with the $o(1)$ error terms set to zero, where we use 20 00020\,000 samples to estimate the population-level statistics $T_{\mathfrak{H}_{1}},T_{\mathfrak{H}_{0}}$ and their standard deviations $\tau_{\mathfrak{H}_{1}},\tau_{\mathfrak{H}_{0}}$. The orange line (empirical) computes the power based on the simulated null distribution, i.e. the threshold estimated via the empirical CDF of $\widehat{T}_{0}$ with test size $m$. The green line (perm) is the permutation test power. All three lines roughly agree with each other.
@@ 3 Tests with the Hilbert-Schmidt Independence Criterion (HSIC)
Suppose we have a kernel $k$ on $\mathcal{X}$ with RKHS $\mathcal{F}$ and feature map $\phi$, as well as another kernel $l$ on $\mathcal{Y}$ with RKHS $\mathcal{G}$ and feature map $\psi$. Let $\otimes$ denote the outer product. [^1] The Euclidean outer product $ab^{\mathsf{T}}$ is a matrix with $ab^{\mathsf{T}}]b^{\prime}=a[b^{\mathsf{T}}b^{\prime}]$; in Hilbert spaces, $f\otimes g:\mathcal{G}\to\mathcal{F}$ has $f\otimes g]g^{\prime}=f\langle g,g^{\prime}\rangle_{\mathcal{G}}$. The cross-covariance operator is
@@ 3.1 Asymptotic test power
The unbiased HSIC estimator is asymptotically normal with $\sqrt{m}(\widehat{\operatorname{HSIC}}_{\rm u}-\operatorname{HSIC})\overset{\mathrm{d}}{\to}\mathcal{N}(0,\sigma^{2}_{\mathfrak{H}_{1}})$ (Proposition˜A.1, in Appendix˜A). It is possible, however, to have $\sigma^{2}_{\mathfrak{H}_{1}}=0$, which makes that result less useful; this occurs in particular whenever $\operatorname{HSIC}=0$, and hence is always true under $\mathfrak{H}_{0}$. In that case, it is more informative to say instead that $m\,\widehat{\operatorname{HSIC}}_{\rm u}$ converges in distribution (which is not the case when $\sigma^{2}_{\mathfrak{H}_{1}}>0$). The distribution to which it converges is a mixture of shifted chi-squareds with complex dependence on $\mathbb{P}_{x}$, $\mathbb{P}_{y}$, $k$, and $l$ (Proposition˜A.2). Thus, if we consider a test statistic $m\,\widehat{\operatorname{HSIC}}_{\rm u}$, when $\operatorname{HSIC}=0$ this statistic has mean zero and standard deviation $\Theta(1)$. When $\operatorname{HSIC}>0$, though, the statistic has mean and standard deviation $\Theta(\sqrt{m})\to\infty$. Thus, as $m\to\infty$, eventually the test will reject if $\operatorname{HSIC}>0$ (Rindt et al., 2021).
@@
Recall that the optimal test is one with highest test power among those with appropriate level control. As for NDS-based tests, we can describe this power asymptotically. Let $\Psi$ be the cdf of the null distribution of $m\,\widehat{\operatorname{HSIC}}_{\rm u}$, which depends on $k$, $l$, $\mathbb{P}_{x}$, and $\mathbb{P}_{y}$. It follows that the rejection threshold $r_{m}$ satisfies
@@
$\alpha=\Pr_{\mathfrak{H}_{0}}(\widehat{\operatorname{HSIC}}_{\rm u}>r_{m})=1-\Psi(mr_{m})+o(1)\qquad\text{so}\qquad r_{m}=\frac{1}{m}\bigl(\Psi^{-1}(1-\alpha)+o(1)\bigr).$
@@
Using Proposition˜A.1, when $\sigma_{\mathfrak{H}_{1}}>0$ the asymptotic test power then satisfies
@@
$\Pr_{\mathfrak{H}_{1}}\left(\widehat{\operatorname{HSIC}}_{\rm u}>r_{m}\right)$ $=\Pr_{\mathfrak{H}_{1}}\left(\sqrt{m}\frac{\widehat{\operatorname{HSIC}}_{\rm u}-\operatorname{HSIC}}{\sigma_{\mathfrak{H}_{1}}}>\sqrt{m}\frac{r_{m}-\operatorname{HSIC}}{\sigma_{\mathfrak{H}_{1}}}\right)$
$=1-\Phi\left(\sqrt{m}\frac{\frac{1}{m}\bigl(\Psi^{-1}(1-\alpha)+o(1)\bigr)-\operatorname{HSIC}}{\sigma_{\mathfrak{H}_{1}}}\right)+o(1)$
$=\Phi\left(\frac{\sqrt{m}\operatorname{HSIC}}{\sigma_{\mathfrak{H}_{1}}}-\frac{\Psi^{-1}(1-\alpha)+o(1)}{\sqrt{m}\,\sigma_{\mathfrak{H}_{1}}}\right)+o(1).$ (6)
@@
Figure˜3 verifies that this expression lines up well with the empirical test power even at moderate test sizes. Additionally, (6) also tells us that, as long as $\operatorname{HSIC}>0$ and $\widehat{\operatorname{HSIC}}_{\rm u}$ has finite positive variance, an HSIC-based test will eventually reject with probability of one. The rate at which it increases in power is described by the gap between the signal-to-noise ratio $\operatorname{HSIC}/\sigma_{\mathfrak{H}_{1}}$ and a threshold-dependent term $\Psi^{-1}(1-\alpha)/\sigma_{\mathfrak{H}_{1}}$, with the former dominating the latter. This gap increases at a faster rate than the one described for NDS-based tests (3); the ratio of the two terms is $\Theta(m)$ for HSIC, while it is only $\Theta(\sqrt{m})$ for NDS. (Appendix˜B gives another motivation for the power being dominated by this signal-to-noise ratio, without using the central limit theorem.)
@@
Figure 3: Power of HSIC tests based on kernels $k$ and $l$ learned with the method of Section˜5, evaluated on problems described in Section˜6. For each problem, the blue line (asymp) is power described by the asymptotic formula (6), where we use 10 00010\,000 samples to estimate the population level statistics $\operatorname{HSIC}$, its deviation $\sigma_{\mathfrak{H}_{1}}$, and the threshold $\Psi^{-1}(1-\alpha)$ via simulation. The orange line (empirical) computes the power based on the simulated null distribution, i.e. the threshold estimated via the empirical CDF of $\widehat{\operatorname{HSIC}}_{u}$ with test size $m$. The green line (perm) is the permutation test power. All three lines roughly agree.
@@ 3.2 MMD-based independence test
Next we consider a testing procedure. The standard independence problem observes paired samples $\mathbb{Z}=\bigl((x_{1},y_{1}),\dots,(x_{m},y_{m})\bigr)$ drawn i.i.d. from the joint distribution $\mathbb{P}_{xy}$, and so to simulate samples from $\mathbb{P}_{x}\times\mathbb{P}_{y}$ we first shuffle the $Y$ samples by some permutation $\sigma$ on $\{1,...,m\}$. We define this action of permuting $\mathbb{Z}$ with $\sigma$ as $\sigma\mathbb{Z}=\bigl((x_{1},y_{\sigma(1)}),\dots,(x_{m},y_{\sigma(m)})\bigr)$. These samples are then used to compute $\widehat{\operatorname{MMD}}_{b}^{2}(\mathbb{Z},\sigma\mathbb{Z})$. The following proposition shows that, with enough samples, this statistic is a consistent estimator of the true $\operatorname{MMD}$ value.
@@
This result is proved in Section˜F.4.
Viewing this permuted MMD estimator as a function of the original paired samples, i.e. $f(\mathbb{Z}):=\widehat{\operatorname{MMD}}_{b}^{2}(\mathbb{Z},\sigma\mathbb{Z})$, we then perform an independence permutation test by shuffling only the $Y$ samples in $\mathbb{Z}$. Under the null hypothesis the distribution of $\mathbb{Z}$ is unchanged by shuffling the $Y$ samples, and so an independence permutation test on $f(\mathbb{Z})$ is guaranteed by Theorem 2 of Hemerik & Goeman (2018) to have correct level control, provided the original sample order is included in the permutations. [^3] One might wonder whether we can instead perform a two-sample permutation test. Consider the test statistic as a function of the pooled samples $f([\mathbb{Z},\sigma\mathbb{Z}])=\widehat{\operatorname{MMD}}_{b}^{2}(\mathbb{Z},\sigma\mathbb{Z})$. The two-sample permutation test involves shuffling the order of $\mathbb{Z}\cup\sigma\mathbb{Z}$, however, exchangeability of $\mathbb{Z}\cup\sigma\mathbb{Z}$ under the null is not guaranteed as variables in the first split are deterministically tied to variables in the second; thus a two-sample permutation test may not be valid. A workaround would be to independently split the data into $\mathbb{Z}_{1}$ and $\mathbb{Z}_{2}$ and then construct the pooled sample as $\mathbb{Z}_{1}\cup\sigma\mathbb{Z}_{2}$, but this process is more involved and potentially lower-power.
It’s also worth noting that, since the MMD estimator is a two-sample statistic, the number of samples of $\mathbb{P}_{xy}$ and $\mathbb{P}_{x}\times\mathbb{P}_{y}$ can be different. As such, we can also consider using multiple permutations $\boldsymbol{\sigma}=\{\sigma_{1},...,\sigma_{p}\}$ rather than just one, and then simulate the null samples as $\boldsymbol{\sigma}\mathbb{Z}=\bigcup\sigma_{i}\mathbb{Z}$. It turns out that if we consider all rotation permutations, $\widehat{\operatorname{MMD}}_{b}^{2}(\mathbb{Z},\boldsymbol{\sigma}\mathbb{Z})$ is exactly the biased HSIC estimator $\widehat{\operatorname{HSIC}}_{b}(\mathbb{X},\mathbb{Y})$. We elaborate on this connection in Appendix F.4 and show that using more permutations improves estimation quality. The HSIC estimator, which effectively uses all permutations, is the minimum variance estimator for this permuted MMD statistic (Proposition˜F.2).